OpenStudy (anonymous):

Alice takes part in a crossword puzzle in a newspaper everyday except Sunday. She can solve on average, 7 of the puzzles. What is the mean and standard deviation she can solve in a week?

OpenStudy (anonymous):

I'm weak at Statistics...Sorry..@hartnn

hartnn (hartnn):

average = mean =7 or am i missing something ?

OpenStudy (calculusfunctions):

There you go, average is mean.

OpenStudy (anonymous):

the questtion wording - Alice solves on average 7 puzzles per week is it ?

hartnn (hartnn):

oh, yes! if it is 7 per day, then answer is different

OpenStudy (anonymous):

Im so sorry. I left some information out. She can solve on average, 7 out of 10 puzzle.

hartnn (hartnn):

then i assume , 10 puzzles asked daily. so on average 7 puzzles solved daily implies 7*6=42 per week and mean per week = 42/ 7 = 6

OpenStudy (anonymous):

i think mean would be 7/10 x 6 = 4.2

OpenStudy (anonymous):

max puzzles she can solve = 6 min puzzles she can solve = 3

hartnn (hartnn):

how u got that??

OpenStudy (anonymous):